# https://vcai.mpi-inf.mpg.de/projects/MetaCap/

### v1

@@ Abstract
Faithful human performance capture and free-view rendering from sparse RGB observations is a long-standing problem in Vision and Graphics. The main challenges are the lack of observations and the inherent ambiguities of the setting, e.g. occlusions and depth ambiguity. As a result, radiance fields, which have shown great promise in capturing high-frequency appearance and geometry details in dense setups, perform poorly when naïvely supervising them on sparse camera views, as the field simply overfits to the sparse-view inputs. To address this, we propose MetaCap, a method for efficient and high-quality geometry recovery and novel view synthesis given very sparse or even a single view of the human. Our key idea is to meta-learn the radiance field weights solely from potentially sparse multi-view videos, which can serve as a prior when fine-tuning them on sparse imagery depicting the human. This prior provides a good network weight initialization, thereby effectively addressing ambiguities in sparse-view capture. Due to the articulated structure of the human body and motion-induced surface deformations, learning such a prior is non-trivial. Therefore, we propose to meta-learn the field weights in a pose-canonicalized space, which reduces the spatial feature range and makes feature learning more effective. Consequently, one can fine-tune our field parameters to quickly generalize to unseen poses, novel illumination conditions as well as novel and sparse (even monocular) camera views. For evaluating our method under different scenarios, we collect a new dataset, WildDynaCap, which contains subjects captured in, both, a dense camera dome and in-the-wild sparse camera rigs, and demonstrate superior results compared to recent state-of-the-art methods on both public and WildDynaCap dataset.
@@ 1 Introduction
How to mitigate the effects of these ambiguities and missing information is the core problem of sparse-view reconstructions and researchers have explored different priors to compensate for the ambiguities of such an ill-posed problem. Earlier works introduce priors on template meshes [24, 30, 85, 23] to solve the above ambiguities. However, meshes suffer from limited resolution and are not easy to incorporate into learning frameworks. Implicit surface representations offer a promising alternative and some methods [60, 95] can learn implicit field priors from large-scale human scan datasets. However, because the amount of data is limited and model capacity is small, they suffer from blurry appearance and poor generalization capability. While many works have recently focused on animatable implicit radiance fields for humans [80, 38, 94, 55, 78, 21], only few  [39, 19, 48, 87] have explored priors for optimizing or fitting a neural implicit field given sparse imagery. They usually only show results for simple objects, but not for complex structures like articulated humans.
@@
Recently,  [69, 72] show meta-learning can help learn powerful priors in neural field methods for few-shot learning. To learn the field prior for personalized high fidelity human capture more effectively, we argue that we should decompose the human model into a coarse-level geometry represented by mesh-based templates [40, 22] and a fine-level neural implicit field [76, 79, 93] accounting for details such as clothing wrinkles. The coarse-level prior from multi-view images is easier to obtain and can serve as a crucial component in fine-level prior learning. At the fine level, we, for the first time in literature, explore learning a meta prior from images for an implicit human representation, which at test time can be optimized using sparse image cues, thereby allowing fast adaptation to unseen poses, novel views, different camera setups, and novel illumination conditions.
@@
Our method, dubbed MetaCap, represents the fine-level human body as a signed distance field (SDF) and an appearance field parameterized with an efficient hashgrid encoding [45], which can be volume rendered [76] into an image. To learn an effective prior for such a representation, we propose to meta-learn the hashgrid parameters solely from multi-view imagery during training in an end-to-end manner, i.e. learning the optimal weights, which, at test time under the sparse settings, yields improved convergence speed and accuracy. However, naïvely meta-learning these implicit parameters, as done in [69, 72] for static scenes, leads to poor performance on complicated settings such as human performance capture (Fig. 5). This is because the human surface is highly articulated and always deforms, thus, learning hashgrid parameters, which live on a spatially fixed grid, is not effective, e.g. the limbs can be mapped onto entirely different hash table parameters depending on their articulation. Therefore, we further propose a space canonicalization step, i.e. we transform points from global space to a canonical pose space. To obtain the transform, we query the nearest transformation of coarse-level human template. We highlight that this canonicalization is not tightly bound to a specific template, but supports deformable human models [22] as well as parametric (piece-wise rigid) body models [40]. Additionally, we introduce occlusion handling, where we use a visibility map to guide the ray sampling, enabling reconstruction of heavily occluded regions at sparse or even monocular inputs at inference time.
@@
We evaluate our method on the DynaCap [22] benchmark, which provides multi-view recordings of humans in a controlled studio setup. However, DynaCap lacks in-the-wild scenes for evaluating the robustness to various environment conditions such as the difference in lighting and camera parameters. Thus, we provide a new dataset, WildDynaCap, comprising of paired dense multi-view studio captures and sparse multi-view in-the-wild recordings for each subject. In summary, our main contributions are:
@@
  • A meta-learning method for high-quality reconstruction of human geometry and appearance from sparse imagery.
@@ 2 Related Work
In the following, we discuss prior works on (1) scene-agnostic implicit representations for novel view synthesis and reconstruction, (2) human-specific sparse-view reconstruction and rendering methods, (3) personalized performance capture and (4) 3D reconstruction using meta-learning. Animatable avatar methods [41, 62, 22, 38, 96, 36, 21, 66], which only take the skeletal pose as input during inference, are not in our scope, since we focus on the reconstruction task, i.e. recovering geometry and appearance from 2D imagery during inference.
Scene-agnostic Implicit Representations. The emergence of implicit representations for 3D scene reconstruction [42, 51, 60, 76] and novel-view synthesis [44, 5, 6, 7] paved the way towards compact and high-fidelity scene representations. Subsequent methods have attempted to improve upon a variety of aspects, such as faster training [45, 16, 9], faster inference [88, 32], sparse [10, 39, 19] and monocular-view [89, 18] reconstruction. In our work, we build upon the scene representation introduced in NeuS2 [79] and Instant-NSR [93]. They leverage multi-grid hash encoding [45] for parameterizing the implicit surface [76] to enable fast and high quality reconstruction. However, they require sufficiently dense multi-view images to ensure accurate geometry and appearance recovery. Due to occlusions and the depth ambiguity inherently present in the sparse setup, simply optimizing such methods with sparse RGB imagery leads to poor reconstructions. More recently, some methods explored different geometry priors [39, 19] or regularizations [87, 48] to enable few-shot novel view synthesis. However, none of them proves to be effective under large camera baselines or sparse $360^{\circ}$ scenarios, where the image observations are very limited and occlusions happen more frequently. We address these challenges by meta-learning an optimal set of initial hashgrid parameters, which effectively serves as a prior during inference on sparse observations, and a human-specific space canonicalization.
@@
Personalized Performance Capture. In contrast to generalizable performance capture methods [60, 95, 78, 50], personalized human capture first builds a person-specific prior on (potentially dense) multi-view data [24, 74, 81, 12, 67, 37] while during inference the avatar can be reconstructed from sparser signals, e.g. a monocular image. Therefore, the goal of such methods is less about cross-identity generalization, but rather highest-fidelity surface and appearance recovery from sparse sensory data while ensuring generalization to different human motions, lighting condition, camera poses, camera numbers, and surface dynamics. Our setting is most closely related to these methods. However, we for the first time in literature propose meta-learning as a prior on the implicit human representation solely from images while experimentally showing that we outperform previous methods operating under the same setting.
Reconstruction with Meta-Learning. Meta-learning is intended to learn from multi tasks or single task [26]. When new observations are presented, domain adaptation or improved performance can be achieved with minimal training iterations. Here, we focus on optimization-based methods [35, 15, 47, 4, 57], especially MAML [15] and Reptile [47]. MetaSDF [69] learns SDF priors and achieves faster inference than auto-decoder methods [51]. Tancik et al. [72] apply a similar strategy to different coordinate-based neural representations, such as image fields and radiance fields [44]. MetaAvatar [77] extends this idea from static to dynamic human representations. After meta-learning on 3D human scans [41], the learned neural SDF weights can be fine-tuned using a few depth maps.
ARAH [78] is methodically the most closely related to our work. However, there are significant differences: 1) They meta-learn human priors from 3D scans while only support image-only supervision during fine-tuning. In stark contrast, our formulation allows meta-learning from potentially sparse images directly as we deeply entangle the volume rendering and hash-grid parameterization with the meta-learning routine. Our end-to-end prior learning and fine-tuning, thus, enables better recovery of high-frequency geometry and appearance. 2) Our whole design, i.e. space canonicalization and scene parameterization, is geared towards efficiency, significantly reducing fine-tuning time, reducing from hours (ARAH) to minutes (MetaCap). 3) We experimentally show that our canonicalization is agnostic to specific template choices, thus, we can also account for loose clothing while ARAH solely shows tightly clothed people.
@@ 3 Method
During the meta-learning phase (Sec. 3.1 and 3.2), we are given a multi-view video using $K_{d}$ cameras, which captures the performer. We denote the images of frame $f$ as $\mathcal{I}_{f}=\{I_{f}^{k}|k\in[1;K_{d}]\}$. We also assume the availability of paired masks $\mathcal{O}_{f}$ and the corresponding skeletal motion $\mathcal{M}_{f}=\{\boldsymbol{\theta}_{f-W},\boldsymbol{\alpha}_{f-W},\boldsymbol{z}_{f-W},...,\boldsymbol{\theta}_{f},\boldsymbol{\alpha}_{f},\boldsymbol{z}_{f}\}$, defined over a window of $W$ frames, with $\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{z}$ representing the joint rotations, the root rotations, and the root translations, respectively. Given this, we aim at meta-learning a personalized and optimal set of weights $\phi^{*}$ for the implicit human representation, effectively serving as a data-driven prior.
[…]
@@ 3.1 Meta-Learning a Reconstruction Prior
where $\mathbf{R}=\{\mathbf{r}_{k}|k\in[1;m]\}$ is a collection of $m$ rays, $\mathcal{L}_{\mathrm{color}}=\frac{1}{m}\sum_{k}\mathcal{H}({\hat{C}(\mathbf{r}_{k})},{C}(\mathbf{r}_{k}))$ is the Huber loss [28], $\mathcal{L}_{\mathrm{eik}}=\frac{1}{mn}\sum_{k,i}\left|1-\langle\mathbf{n}(\mathbf{p}_{k,i}),\left.\nabla_{\mathbf{p}_{k,i}}f_{\phi}(\mathbf{p}_{k,i})\right\rangle\right|_{2}^{2}$ is the Eikonal loss [17] on the normal $\mathbf{n}(\mathbf{p}_{k,i})$ of the sample point $\mathbf{p}_{k,i}$. $\mathcal{L}_{\mathrm{mask}}=\frac{1}{m}\sum_{k}BCE(\hat{M}{(\mathbf{r}_{k})},{O}{(\mathbf{r}_{k})})$ is the binary cross entropy loss. $s(\mathbf{p}_{k,i})$ is the SDF value of point $\mathbf{p}_{k,i}$, $\mathcal{L}_{\mathrm{sparse}}=\frac{1}{mn}\sum_{k,i}\exp^{-|s(\mathbf{p}_{k,i})|}$ is a sparseness regularization term [39]. ${C}(\mathbf{r}_{k})$ and ${O}(\mathbf{r}_{k})$ are the ground truth color and mask values of ray $\mathbf{r}_{k}$. To accelerate the rendering process, we parameterize the implicit field using a multi-resolution hashgrid [45]. Thus, the overall learnable parameters $\phi$ include the MLP weights of the implicit function as well as the hashgrid parameters.
@@
where $T_{i}$ are the tasks sampled at the $i^{th}$ outer iteration, $l_{\mathrm{out}}$ denotes the learning rate of outer loop. $\phi_{M}^{\prime}(\phi_{i-1},T_{i})$ represents the network weights initialized by the inner-loop with $\phi_{i-1}$, which are then optimized for $M$ steps on $T_{i}$. The resulting weights, $\phi_{N}$, at the end of $N$ outer loop iterations serve as the optimal initial weights $\phi^{*}$ for the fine-tuning stage. To keep the meta-learning computation tractable, we choose Reptile [47] in Eq. 4 as it uses a first order approximation instead of MAML’s second order gradient update.
@@ 3.2 Template-guided Meta-Learning
Human Template. Here, we introduce a generic motion-driven human template, defined as a deformable mesh $\hat{\mathbf{X}}\in\mathbb{R}^{V\times 3}$, of a character in the canonical pose (as shown in Fig. 2). The template mesh $\hat{\mathbf{X}}$ can be deformed according to a given skeletal motion $\mathcal{M}_{f}$. Each vertex in the human template first undergoes a per-vertex motion-dependent deformation in the canonical space through a transform $\mathbf{T}_{\mathrm{def}}(\mathcal{M}_{f})\in\mathbb{R}^{4\times 4}$. After being deformed locally with $\mathbf{T}_{\mathrm{def}}$, one can perform forward kinematics $\mathbf{T}_{\mathrm{FK}}(\mathcal{M}_{f})$ to transform the deformed canonical space vertices $\mathbf{T}_{\mathrm{def}}\hat{\mathbf{X}}$ to world space vertices ${\mathbf{X}}$. The total invertible transformation from the template’s canonical space to world space is: $\mathbf{T}=\mathbf{T}_{\mathrm{FK}}\mathbf{T}_{\mathrm{def}}$ and the template vertices are transformed as $\mathbf{X}=\mathbf{T}\hat{\mathbf{X}}$.
Template-guided Ray Warping. The template mesh defined above provides us a coarse geometry information, which effectively guides our meta-learning since we are now able to learn the meta-weights in a pose-canonical space. The sampled points $\mathbf{p}_{i}$ can be warped from the world space into the canonical space through our proposed template guided ray warping. Concretely, we first project each sampled point onto the nearest face (triangle) of the posed template mesh $\mathbf{X}$. To compute the transformation matrix $\mathbf{T}_{i}$, we barycentricaly interpolate the transformation of the vertices defining the nearest triangle. Next, we apply the inverse transformation $\mathbf{T}_{i}^{-1}$ of the sampled point to obtain its canonical position $\hat{\mathbf{p}}_{i}=\mathbf{T}_{i}^{-1}\mathbf{p}_{i}$. The canonical points, $\hat{\mathbf{p}_{i}}$, can now be encoded with the learnable hash encoding and rendered using the volumetric rendering approach described in Eq. 1. Moreover the template allows us to perform empty skipping [25] for point samples whose distance to the template exceeds a threshold $\eta$, thereby significantly accelerating the convergence. In this way, our final meta-learning approach (see also Alg. 1) is more efficient and effective compared to the baseline, i.e. learning in global space.
@@
         Transform points into canonical space $\hat{\mathbf{p}}=\mathbf{T}^{-1}\mathbf{p}$
         $\phi^{{}^{\prime}}_{j}=\phi^{{}^{\prime}}_{j-1}-l^{in}\nabla_{\phi}\mathcal{L}(f_{\phi}(\hat{\mathbf{R}}^{(j)})|_{\phi=\phi^{{}^{\prime}}_{j-1}})$

### v2

@@ Abstract
Faithful human performance capture and free-view rendering from sparse RGB observations is a long-standing problem in Vision and Graphics. The main challenges are the lack of observations and the inherent ambiguities of the setting, e.g. occlusions and depth ambiguity. As a result, radiance fields, which have shown great promise in capturing high-frequency appearance and geometry details in dense setups, perform poorly when naïvely supervising them on sparse camera views, as the field simply overfits to the sparse-view inputs. To address this, we propose MetaCap, a method for efficient and high-quality geometry recovery and novel view synthesis given very sparse or even a single view of the human. Our key idea is to meta-learn the radiance field weights solely from potentially sparse multi-view videos, which can serve as a prior when fine-tuning them on sparse imagery depicting the human. This prior provides a good network weight initialization, thereby effectively addressing ambiguities in sparse-view capture. Due to the articulated structure of the human body and motion-induced surface deformations, learning such a prior is non-trivial. Therefore, we propose to meta-learn the field weights in a pose-canonicalized space, which reduces the spatial feature range and makes feature learning more effective. Consequently, one can fine-tune our field parameters to quickly generalize to unseen poses, novel illumination conditions as well as novel and sparse (even monocular) camera views. For evaluating our method under different scenarios, we collect a new dataset, WildDynaCap, which contains subjects captured in, both, a dense camera dome and in-the-wild sparse camera rigs, and demonstrate superior results compared to recent state-of-the-art methods on, both, public and WildDynaCap dataset.
@@ 1 Introduction
How to mitigate the effects of these ambiguities and missing information is the core problem of sparse-view reconstructions and researchers have explored different priors to compensate for the ambiguities of such an ill-posed problem. Earlier works introduce priors on template meshes [24, 30, 91, 23] to solve the above ambiguities. However, meshes suffer from limited resolution and are not easy to incorporate into learning frameworks. Implicit surface representations offer a promising alternative and some methods [65, 101] can learn implicit field priors from large-scale human scan datasets. However, they suffer from blurry appearance and poor generalization capability because the amount of data is limited and model capacity is small. While many works have recently focused on animatable implicit radiance fields for humans [86, 41, 100, 60, 84, 21], only few  [42, 19, 52, 93] have explored priors for optimizing or fitting a neural implicit field given sparse imagery. They usually only show results for simple objects, but not for complex structures like articulated humans.
@@
Recently, some works [74, 78] showed that meta-learning can learn powerful neural field priors for few-shot learning. To learn the field prior for personalized high fidelity human capture more effectively, we argue that we should decompose the human model into a coarse-level geometry represented by mesh-based templates [43, 22] and a fine-level neural implicit field [82, 85, 99] accounting for details such as clothing wrinkles. The coarse-level prior from multi-view images is easier to obtain and can serve as a crucial component in fine-level prior learning. At the fine level, we, for the first time in literature, explore learning a meta prior from images for an implicit human representation, which at test time can be optimized using sparse image cues, thereby allowing fast adaptation to unseen poses, novel views, different camera setups, and novel illumination conditions.
@@
Our method, dubbed MetaCap, represents the fine-level human body as a signed distance field (SDF) and an appearance field parameterized with an efficient hashgrid encoding [49], which can be volume rendered [82] into an image. To learn an effective prior for such a representation, we propose to meta-learn the hashgrid parameters solely from multi-view imagery during training in an end-to-end manner, i.e. learning the optimal weights, which, at test time under the sparse settings, yield improved convergence rate and accuracy. However, naïvely meta-learning these implicit parameters for static scenes [74, 78], leads to poor performance on complicated settings such as human performance capture (Fig. 5). This is because the human surface is highly articulated and always deforms. Thus, learning hashgrid parameters, which live on a spatially fixed grid, is not effective, e.g. the limbs can be mapped onto entirely different hash table parameters depending on their articulation. Therefore, we further propose a space canonicalization step, i.e. we transform points from global space to a canonical pose space. To obtain the transform, we query the nearest transformation of a coarse-level human template. We highlight that this canonicalization is not tightly bound to a specific template, but supports deformable human models [22] as well as parametric (piece-wise rigid) body models [43]. Additionally, we introduce occlusion handling, where we use a visibility map to guide the ray sampling, enabling reconstruction of heavily occluded regions at sparse or even monocular inputs at inference time.
@@
We evaluate our method on the DynaCap [22] benchmark, which provides multi-view recordings of humans in a controlled studio setup. However, DynaCap lacks in-the-wild scenes for evaluating the robustness to various environment conditions such as the difference in lighting and camera parameters. Thus, we provide a new dataset, WildDynaCap, containing paired dense multi-view studio captures and sparse multi-view in-the-wild recordings for each subject. In summary, our main contributions are:
@@
  • A meta-learning method learned from multi-view imagery for high-quality reconstruction of human geometry and appearance under sparse cameras.
@@ 2 Related Work
In the following, we discuss prior works on (1) scene-agnostic implicit representations for novel view synthesis and reconstruction, (2) human-specific sparse-view reconstruction and rendering methods, (3) personalized performance capture and (4) 3D reconstruction using meta-learning. Animatable avatar methods [45, 67, 22, 41, 102, 38, 21, 71, 36, 55, 44, 105], which only take the skeletal pose as input during inference, are not in our scope, since we focus on the reconstruction task, i.e. recovering geometry and appearance from 2D imagery during inference.
Scene-agnostic Implicit Representations. The emergence of implicit representations for 3D scene reconstruction [46, 56, 65, 82] and novel-view synthesis [48, 5, 6, 7] paved the way towards compact and high-fidelity scene representations. Subsequent methods have attempted to improve upon a variety of aspects, such as faster training [49, 16, 9], faster inference [94, 33], sparse [10, 42, 19] and monocular-view [95, 18] reconstruction. In our work, we build upon the scene representation introduced in NeuS2 [85] and Instant-NSR [99]. They leverage multi-grid hash encoding [49] for parameterizing the implicit surface [82] to enable fast and high quality reconstruction. However, they require sufficiently dense multi-view images to ensure accurate geometry and appearance recovery. Due to occlusions and the depth ambiguity inherently present in the sparse setup, simply optimizing such methods with sparse RGB imagery leads to poor reconstructions. More recently, some methods explored different geometry priors [42, 19] or regularizations [93, 52, 31] to enable few-shot novel view synthesis. However, none of them proves to be effective under large camera baselines or sparse $360^{\circ}$ scenarios, where the image observations are very limited and occlusions happen more frequently. We address these challenges by meta-learning an optimal set of initial hashgrid parameters, which effectively serves as a prior during inference on sparse observations, and a human-specific space canonicalization.
@@
Personalized Performance Capture. In contrast to generalizable performance capture methods [65, 101, 84, 54], personalized human capture first builds a person-specific prior on multi-view data [24, 80, 87, 12, 72, 40, 39] while during inference the avatar can be reconstructed from sparser signals, e.g. a monocular image. Therefore, the goal of such methods is less about cross-identity generalization, but rather highest-fidelity surface and appearance recovery from sparse sensory data while ensuring generalization to different human motions, lighting condition, camera poses, camera numbers, and surface dynamics. Our setting is most closely related to these methods. However, we for the first time in literature propose meta-learning as a prior on the implicit human representation solely from images while experimentally showing that we outperform previous methods operating under the same setting.
Reconstruction with Meta-Learning. Meta-learning is intended to learn from multiple tasks or a single task [26]. When new observations are presented, domain adaptation or improved performance can be achieved with minimal training iterations. Here, we focus on optimization-based methods [37, 15, 51, 4, 62], especially MAML [15] and Reptile [51]. MetaSDF [74] learns SDF priors and achieves faster inference than auto-decoder methods [56]. Tancik et al. [78] apply a similar strategy to different coordinate-based neural representations, such as image fields and radiance fields [48]. MetaAvatar [83] extends this idea from static to dynamic human representations. After meta-learning on 3D human scans [45], the learned SDF weights can be fine-tuned using a few depth maps.
ARAH [84] is methodically most closely related to our work. However, there are significant differences: 1) They meta-learn human priors from 3D scans while only supporting image-only supervision during fine-tuning. In stark contrast, our formulation allows meta-learning from potentially sparse images directly as we deeply entangle the volume rendering and hash-grid parameterization with the meta-learning routine. Our end-to-end prior learning and fine-tuning, thus, enables better recovery of high-frequency geometry and appearance. 2) Our whole design, i.e. space canonicalization and scene parameterization, is geared towards efficiency, significantly improving the fine-tuning time, reducing from hours (ARAH) to minutes (MetaCap). 3) We experimentally show that our canonicalization is agnostic to specific template choices, thus, we can also account for loose clothing while ARAH solely shows tightly clothed people.
@@ 3 Method
During the meta-learning phase (Sec. 3.1 and 3.2), we are given a multi-view video using $K_{d}$ cameras, which captures the performer. We denote the images of frame $f$ as $\mathcal{I}_{f}=\{I_{f}^{k}|k\in[1;K_{d}]\}$ and assume the availability of paired masks $\mathcal{O}_{f}$ and the corresponding skeletal motion $\mathcal{M}_{f,W}=\{\boldsymbol{\theta}_{f-W},\boldsymbol{\alpha}_{f-W},\boldsymbol{z}_{f-W},...,\boldsymbol{\theta}_{f},\boldsymbol{\alpha}_{f},\boldsymbol{z}_{f}\}$, defined over a window of $W$ frames, with $\boldsymbol{\theta},\boldsymbol{\alpha},\boldsymbol{z}$ representing the joint rotations, the root rotation, and the root translation, respectively. Given this, we aim at meta-learning a personalized and optimal set of weights $\phi^{*}$ for the implicit human representation, effectively serving as a data-driven prior.
During the fine-tuning phase (Sec. 3.3), we fine-tune the meta-learned weights $\phi^{*}$ to sparse in-the-wild images $\mathcal{{I}}_{f}=\{{I}_{f}^{k}|k\in[1;K_{s}]\}$ paired with masks ${\mathcal{O}}_{f}=\{{O}_{f}^{k}|k\in[1;K_{s}]\}$ with $K_{s}<<K_{d}$. Notably, the motion $\mathcal{{M}}_{f,W}$ and surface dynamics of the performer as well as the lighting conditions and camera configuration during fine-tuning can significantly differ from the ones observed during training. We later show that fine-tuning the meta-learned weights $\phi^{*}$ on $\mathcal{{I}}_{f}$ leads to faster and better convergence compared to random weights $\phi_{0}$.
@@ 3.1 Meta-Learning a Reconstruction Prior
where $\mathbf{R}=\{\mathbf{r}_{k}|k\in[1;m]\}$ is a collection of $m$ rays, $\mathcal{L}_{\mathrm{color}}=\frac{1}{m}\sum_{k}\mathcal{H}({\hat{C}(\mathbf{r}_{k})},{C}(\mathbf{r}_{k}))$ is the Huber loss [28], $\mathcal{L}_{\mathrm{eik}}=\frac{1}{mn}\sum_{k,i}\left|1-\langle\mathbf{n}(\mathbf{p}_{k,i}),\left.\nabla_{\mathbf{p}_{k,i}}f_{\phi}(\mathbf{p}_{k,i})\right\rangle\right|_{2}^{2}$ is the Eikonal loss [17] on the normal $\mathbf{n}(\mathbf{p}_{k,i})$ of the sample point $\mathbf{p}_{k,i}$. $\mathcal{L}_{\mathrm{mask}}=\frac{1}{m}\sum_{k}BCE(\hat{M}{(\mathbf{r}_{k})},{O}{(\mathbf{r}_{k})})$ is the binary cross entropy loss. $s(\mathbf{p}_{k,i})$ is the SDF value of point $\mathbf{p}_{k,i}$, and $\mathcal{L}_{\mathrm{sparse}}=\frac{1}{mn}\sum_{k,i}\exp^{-|s(\mathbf{p}_{k,i})|}$ is a sparseness regularization term [42]. ${C}(\mathbf{r}_{k})$ and ${O}(\mathbf{r}_{k})$ are the ground truth color and mask values of ray $\mathbf{r}_{k}$. To accelerate the rendering process, we parameterize the implicit field using a multi-resolution hashgrid [49]. Thus, the overall learnable parameters $\phi$ include the MLP weights of the implicit function as well as the hashgrid parameters.
@@
where $T_{i}$ are the tasks sampled at the $i^{th}$ outer iteration, $l_{\mathrm{out}}$ denotes the learning rate of outer loop. $\phi_{M}^{\prime}(\phi_{i-1},T_{i})$ represents the network weights initialized by the inner-loop with $\phi_{i-1}$, which are then optimized for $M$ steps on $T_{i}$. The resulting weights, $\phi_{N}$, at the end of $N$ outer loop iterations serve as the optimal initial weights $\phi^{*}$ for the fine-tuning stage. To keep the meta-learning computation tractable, we choose Reptile [51] in Eq. 4 as it uses a first-order approximation instead of MAML’s second order gradient update.
@@ 3.2 Template-guided Meta-Learning
Human Template. Here, we introduce a generic motion-driven human template, defined as a deformable mesh $\bar{\mathbf{X}}\in\mathbb{R}^{V\times 3}$, of a character in the canonical pose $\bar{\mathcal{M}}$ (as shown in Fig. 2). The template mesh $\bar{\mathbf{X}}$ can be deformed according to a given skeletal motion $\mathcal{M}_{f,W}$. Each vertex in the human template first undergoes a per-vertex motion-dependent deformation in the canonical space through a transform $\mathbf{T}_{\mathrm{def}}(\mathcal{M}_{f,W})\in\mathbb{R}^{4\times 4}$. After being deformed locally with $\mathbf{T}_{\mathrm{def}}$, one can perform forward kinematics $\mathbf{T}_{\mathrm{FK}}(\mathcal{M}_{f},\bar{\mathcal{M}})$ to transform the deformed canonical space vertices $\mathbf{T}_{\mathrm{def}}\bar{\mathbf{X}}$ to world space vertices ${\mathbf{X}}$. The total invertible transformation from the template’s canonical space to world space is: $\mathbf{T}=\mathbf{T}_{\mathrm{FK}}\mathbf{T}_{\mathrm{def}}$ and the template vertices are transformed as $\mathbf{X}=\mathbf{T}\bar{\mathbf{X}}$.
Template-guided Ray Warping. The template mesh defined above provides us a coarse geometry information, which effectively guides our meta-learning since we are now able to learn the meta-weights in a pose-canonical space. The sampled points $\mathbf{p}_{i}$ can be warped from the world space into the canonical space through our proposed template guided ray warping. Concretely, we first project each sampled point onto the nearest face (triangle) of the posed template mesh $\mathbf{X}$. To compute the transformation matrix $\mathbf{T}_{i}$, we barycentricaly interpolate the transformation of the vertices defining the nearest triangle. Next, we apply the inverse transformation $\mathbf{T}_{i}^{-1}$ of the sampled point to obtain its canonical position $\bar{\mathbf{p}}_{i}=\mathbf{T}_{i}^{-1}\mathbf{p}_{i}$. The canonical points, $\bar{\mathbf{p}_{i}}$, can now be encoded with the learnable hash encoding and rendered using the volumetric rendering approach described in Eq. 1. Moreover the template allows us to perform empty skipping [25] for point samples whose distance to the template exceeds a threshold $\eta$, thereby significantly accelerating the convergence. In this way, our final meta-learning approach (see also Alg. 1) is more efficient and effective compared to the baseline, i.e. learning in global space.
@@
         Transform points into canonical space $\bar{\mathbf{p}}=\mathbf{T}^{-1}\mathbf{p}$
         $\phi^{{}^{\prime}}_{j}=\phi^{{}^{\prime}}_{j-1}-l^{in}\nabla_{\phi}\mathcal{L}(f_{\phi}(\bar{\mathbf{R}}^{(j)})|_{\phi=\phi^{{}^{\prime}}_{j-1}})$